my mom was 27 when i was born.
8 years ago she was twice as old as i shall be in 5 years time.
how old am i now?

Answer by richwmiller(17219)

It is nice that you know all about your mom's age but very sad that you don't know how old you are. You should talk to her and have her tell you stories of your childhood. Do you know your name and where you are? You might have amnesia!
m=c+27,
m-8=2*(c+5)
You are only 9 years old and your mom is 36.
I hope you can talk to your dad about your problems.